# Incremental Joint Learning of Depth, Pose and Implicit Scene Representation on Monocular Camera in Large-scale Scenes

### v1

@@ Abstract
Dense scene reconstruction for photo-realistic view synthesis has various applications, such as VR/AR, autonomous vehicles. However, most existing methods have difficulties in large-scale scenes due to three core challenges: (a) inaccurate depth input. Accurate depth input is impossible to get in real-world large-scale scenes. (b) inaccurate pose estimation. Most existing approaches rely on accurate pre-estimated camera poses. (c) insufficient scene representation capability. A single global radiance field lacks the capacity to effectively scale to large-scale scenes. To this end, we propose an incremental joint learning framework, which can achieve accurate depth, pose estimation, and large-scale scene reconstruction. A vision transformer-based network is adopted as the backbone to enhance performance in scale information estimation. For pose estimation, a feature-metric bundle adjustment (FBA) method is designed for accurate and robust camera tracking in large-scale scenes. In terms of implicit scene representation, we propose an incremental scene representation method to construct the entire large-scale scene as multiple local radiance fields to enhance the scalability of 3D scene representation. Extended experiments have been conducted to demonstrate the effectiveness and accuracy of our method in depth estimation, pose estimation, and large-scale scene reconstruction.
@@ I Introduction
In the past two decades, the demand for dense scene reconstruction has increased across various applications, including mobile robotics [1, 2], autonomous driving [3, 4, 5, 6], smart warehouse [7] and virtual reality [8, 9, 10]. Many works focused on this field have made significant progress using Neural Radiance Fields (NeRF) [11]. The original NeRF method needs the pre-compute pose input from COLMAP or SLAM methods [12, 13, 14, 15], which makes it challenging to use in real-world deployments due to the long processing time. Structure from motion (SFM) is also not robust and accurate in hand-held scenes. Some recent works are proposed to simultaneously estimate camera poses and neural implicit representation, such as BARF [16], SC-NeRF [17], and so on [18, 19, 20]. However, existing methods come with the assumptions of a fixed-size map, which renders them useless in the real-world application.
Figure 1: Qualitative results on the Tanks and Template datasets [21] on Ballroom sequence. Our method renders better novel view synthesis than existing methods.
Recent works, such as NoPe-NeRF [22] and LocalRF [23], have been proposed to improve the scene representation, which is closely related to our work. They design a joint learning approach for pose and scene representation. NoPe-NeRF also estimates the scale correction parameters during training. LocalRF decomposes the scene into several local radiance fields and progressively estimates the poses of input frames. However, both of them have difficulty in dealing with long trajectories and accumulating pose drifts. They are also sensitive to pixel-wise noise variation in longer trajectory sequences.
@@
To this end, we propose an incremental joint learning method for depth, pose estimation, and scene implicit representation. The depth estimation network adopts a vision transformer-based network to encode input images into high-level features. The transformer backbone utilizes a global receptive field to process features at every stage, which provides a more precise estimation. For the pose estimation network, a feature-metric bundle adjustment (FBA) method is proposed for robust pose estimation. Compared to the existing BA methods, our FBA method effectively avoids the sensitivity of pixel-wise noise variation, enabling robust localization in large-scale scenes. For scene implicit representation, existing methods usually use an MLP to encode the scene, leading to poor scene representation capability. We incorporate the feature grid-based NeRF with an incremental scene representation method. We dynamically initialize local radiance fields when the camera moves to the bound of the local scene representation. The incremental scene representation method can significantly enhance the scene representation capability in large-scale scenes.
@@ II-A Joint Learning of Depth and Pose
Some early works [24] use groundtruth depth maps as supervision to learn monocular depth estimation from monocular images. [25] use synthetic images to help the disparity learning. Recently, some works have focused on self-supervised learning [26, 27, 28, 29] to jointly estimate pose and depth.  [30] and [31] jointly optimize pose and depth via photometric reprojection error. [32] focus on the scale ambiguity for monocular depth estimation with normalization. However, convolutional neural networks do not have a global receptive field at every stage and fail to capture global consistency.
@@ II-B Implicit Scene Representation
With the proposal of NeRF [11], many researchers explore combining this implicit method into various applications,such as autonomous driving [33] and object detection [34, 35, 36]. However, the original NeRF method needs groundtruth pose information as their input, which is difficult to obtain in real-world environments. They require pre-computed camera parameters obtained from SFM and SLAM algorithms [37, 38, 39]. Recently, many researchers have focused on implicit scene representation without pose prior. iNeRF [40] is the first work focused on pose estimation and scene representation. They estimate the camera pose for novel view images with a pre-trained NeRF model. BARF [16] proposes a BA optimization for camera poses with and NeRF. Some other methods [41, 42, 43, 44] combine semantic feature with NeRF to improve the scene representation and camera tracking. NoPe-NeRF [22] and LocalRF [23] are the most relevant work to ours. They jointly estimate depth, pose, and scene representation. However, they often fail in large-scale scene representation, and their pose estimation module could be more robust.
@@ III Joint learning of depth, pose, and implicit scene representation
Figure 2: The pipeline of our system. There are three modules in our system: depth estimation module, FBA pose estimation module, and incremental scene implicit representation module. All three modules are jointly optimized during the system operation.
@@ III-A Depth Estimation
Tanks and Temples [21] Ballroom 20.66 0.50 0.60 24.97 0.71 0.39 25.29 0.89 0.08 26.50 0.85 0.08
Barn 25.28 0.64 0.48 25.77 0.68 0.44 27.89 0.88 0.11 28.59 0.91 0.11
Church 23.17 0.62 0.52 23.60 0.67 0.46 28.07 0.91 0.07 28.82 0.92 0.08
Family 23.04 0.61 0.56 23.77 0.68 0.48 28.60 0.92 0.007 30.48 0.94 0.05
Francis 25.85 0.69 0.57 29.48 0.80 0.38 31.71 0.93 0.11 33.38 0.93 0.10
Horse 24.09 0.72 0.41 27.30 0.83 0.27 29.98 0.94 0.07 30.07 0.93 0.07
Ignatius 21.78 0.47 0.60 23.77 0.61 0.47 26.54 0.86 0.09 27.33 0.87 0.09
Museum 23.58 0.61 0.55 26.39 0.76 0.36 26.35 0.86 0.16 29.43 0.92 0.12
Mean 23.42 0.61 0.54 26.34 0.74 0.39 27.87 0.91 0.09 29.04 0.90 0.09
III-B Feature-metric BA Pose Estimation
In terms of pose estimation, we propose a FBA network. We show the architecture of our network in Fig. 2. Inspired by [48], each image is encoded into high-level features and uses FBA to get an accurate and robust pose estimation. For every consecutive frames $I_{a},I_{b}$, we use a Unet-like encoder-decoder network to extract $D_{l}$ dimensional feature map $\mathbf{F}^{l}\in\mathbb{R}^{W_{l}\times H_{l}\times D_{l}}$ at each level $l\in\{L,\ldots,1\}$. Those multi-channel image features with decreasing resolution can encode richer semantic information and larger spatial information of the image. Our network can also simultaneously estimate the confidence of each pixel to select the robust features for pose estimation.
Compared with the original BA method, this learned representation is robust to large illumination or viewpoint changes, which can provide accurate pose estimation in indoor and outdoor environments. Despite poor initial pose estimates, it can also provide meaningful gradients for successful estimation. Those joint learning methods with classical direct alignment [22, 23] operate on the original RGB image, which is not robust to large-scale scenes and long-term changes, and resorts to Gaussian image pyramids, which still largely limits the convergence to frame-to-frame tracking.
Feature Residual: We define the direct feature residual between the two consecutive frames. We optimize the relative pose$\{R,t\}$ through minimizing the feature residual. For a given feature level $l$ and each 3D point $i$ observed in two consecutive frames, we define a residual:
where $\Pi\left(\mathbf{R}\mathbf{P}_{i}+\mathbf{t}\right)_{b}$ is the projection of frame b given its current pose estimate. [·] is a lookup with sub-pixel interpolation. $F_{a}^{l},F_{b}^{l}$ are the corresponding feature maps of images a and b. The total error over N observations is:
Figure 4: Qualitative results on Static Hikes Indoor datasets [23] with 40m trajectory and span over 1000 images. We present the scene reconstruction results of NoPe-NeRF[22], LocalRF [23], and our method in different time steps (90, 220, 390). Our incremental scene representation method significantly improves the scene reconstruction performance in large-scale scenes.
@@ III-C Incremental Implicit Scene Representation
Existing scene implicit scene representation methods such as [17, 49, 16] can only achieve satisfactory results in small-scale scenes. They are frequently constrained by their limitations in scene representation capabilities when dealing with longer sequences and tend to converge towards local minima during the optimization process. Inspired by [23], we propose an incremental implicit scene representation. The architecture of our network is shown in Fig. 2. At the beginning, we will initialize a local neural radiance field and optimize it using the initial few frames as input. As the robot moves forward, we progressively append new frames within the local neural radiance field and iteratively refine the scene representation, pose estimation, and depth estimation. This procedure persists until we reach the boundary of the local neural radiance field. Then, we dynamically allocate new local radiance fields throughout the optimization with frames within a sliding window. This procedure will continue until all the images are registered into the network. So, the entire scene can be represented as multiple local NeRF:
@@
$T_{i}$ $=\exp\left(-\sum_{i-1}^{j}\sigma_{j}\delta_{j}\right)$
III-D Incremental Joint Learning
$\mathcal{L}_{d}=\left|\hat{\mathbf{D}}^{*}-\mathbf{D}^{*}\right|$ (10)
where $\hat{D}^{*}(r)$ is our render depth, $D^{*}(r)$ is our estimated depth image from our depth network. We use a pre-trained model to get depth images prior to our depth network. We define a photo-metric loss to optimize local radiance fields:
@@ III-D Incremental Joint Learning
This is the forward optical flow and we also calculate backward optical flow $L_{fb}$ to supervise our implicit scene representation network. The final loss of our network is :
@@
TABLE III: Novel view synthesis on Static Hikes dataset [23] and Proprietary dataset. We use PSNR, SSIM, and LPIPS as our metrics.
Methods Static Hikes Indoor [23] Proprietary Dataset

### v2

@@ Abstract
Dense scene reconstruction for photo-realistic view synthesis has various applications, such as VR/AR, autonomous vehicles. However, most existing methods have difficulties in large-scale scenes due to three core challenges: (a) inaccurate depth input. Accurate depth input is impossible to get in real-world large-scale scenes. (b) inaccurate pose estimation. Existing methods are not robust enough with the growth of cumulative errors in large scenes and long sequence. (c) insufficient scene representation capability. A single global radiance field lacks the capacity to effectively scale to large-scale scenes. To this end, we propose an incremental joint learning framework, which can achieve accurate depth, pose estimation, and large-scale scene reconstruction. A vision transformer-based network is adopted as the backbone to enhance performance in scale information estimation. For pose estimation, a feature-metric bundle adjustment (FBA) method is designed for accurate and robust camera tracking in large-scale scenes. In terms of implicit scene representation, we propose an incremental scene representation method to construct the entire large-scale scene as multiple local radiance fields to enhance the scalability of 3D scene representation. Extended experiments have been conducted to demonstrate the effectiveness and accuracy of our method in depth estimation, pose estimation, and large-scale scene reconstruction. The code will be open-sourced upon paper acceptance.
Index Terms:
Neural Radiance Fields, Incremental Joint Learning, Pose Estimation, Large-scale Scenes.
@@ I Introduction
In the past two decades, the demand for dense scene reconstruction has increased across various applications, including mobile robotics, autonomous driving. Many works focused on this field have made significant progress using Neural Radiance Fields (NeRF) [1]. The original NeRF method needs the pre-compute pose input from COLMAP or SLAM methods [2], which makes it challenging to use in real-world deployments due to the long processing time. Structure from motion (SFM) is also not robust and accurate in hand-held scenes. Some recent works are proposed to simultaneously estimate camera poses and neural implicit representation, such as BARF [3], SC-NeRF [4], and so on [5, 6, 7]. However, existing methods come with the assumptions of a fixed-size map, which renders them useless in the real-world application.
Figure 1: Qualitative results on the Tanks and Template datasets [8] on Ballroom sequence. Our method achieve better view synthesis performance than existing methods.
Recent works, such as NoPe-NeRF [9] and LocalRF [10], have been proposed to improve the scene representation, which is closely related to our work. They design a joint learning approach for pose and scene representation. NoPe-NeRF also estimates the scale correction parameters during training. LocalRF progressively estimates the poses of input frames. However, both of them have difficulty in dealing with long trajectories and accumulating pose drifts. Their end-to-end pose estimation are also sensitive to pixel-wise noise variation in longer trajectory sequences.
@@
To this end, we propose an incremental joint learning method for depth, pose estimation, and scene implicit representation. These three elements are closely interconnected: depth information is crucial for geometric representation of the scene. It can guide the neural point sampling and enhance scene geometry consistency. Accurate pose estimation is also vital for scene representation. So, we design a joint learning framework for these three elements. The depth estimation network adopts a vision transformer-based network to encode input images into high-level features. The transformer backbone utilizes a global receptive field to process features at every stage, which provides a more precise estimation. For the pose estimation network, a feature-metric bundle adjustment (FBA) method is proposed for robust pose estimation. Compared to the existing BA methods, our FBA method effectively avoids the sensitivity of pixel-wise noise variation, enabling robust localization in large-scale scenes. For scene implicit representation, existing methods usually use an MLP to encode the scene, leading to poor scene representation capability. We propose an incremental scene representation method. We dynamically initialize local radiance fields when the camera moves to the bound of the local scene representation. The incremental scene representation method can significantly enhance the scene representation capability in large-scale scenes.
@@ II-A Joint Learning of Depth and Pose
Some early works [11, 12] use groundtruth depth maps as supervision to learn monocular depth estimation from monocular images. [13] use synthetic images to help the disparity learning. Recently, some works have focused on self-supervised learning to jointly estimate pose and depth.  [14, 15, 16, 17, 18, 19] and [20] jointly optimize pose and depth via photometric reprojection error. [21] focus on the scale ambiguity for monocular depth estimation with normalization. [22] proposes a masked GAN network for unsupervised monocular depth and ego-motion estimation. [23] jointly perform depth and camera pose estimation and semantic scene segmentation. However, convolutional neural networks do not have a global receptive field at every stage and fail to capture global consistency. [24] incorporate channelwise attention mechanism spatialwise attention mechanism for accurate depth estimation.
@@ II-B Implicit Scene Representation
Early Novel View Synthesis methods applied interpolations between pixels [25], later works use meshes, Multi-Plane Images, and light field [26]. With the proposal of NeRF [1], many researchers explore combining this implicit method into various applications, such as autonomous driving [27] and robot navigation. [28] integrates neural radiance fields with navigation networks end-to-end as a memory structure for the first time, highlighting the potential applications of neural radiance fields in downstream control domain tasks. [29] design a waypoint-based task decomposition strategy which effectively enhances the reliability of learning-based navigation. The key idea of NeRF is to model the radiance field of a scene with a multi-layer perceptron (MLP). However, the representation of volume densities can not commit the geometric consistency, leading to poor surface prediction for reconstruction tasks. UNISURF [30] and NeuS [31] are proposed to combine world-centric 3D geometry representation with neural radiance fields. They replace the volume density with Signed Distance Field (SDF) values, which can further improve the surface reconstruction. Other methods [32, 33] use various scene geometry representation methods, such as truncated signed distance function, voxel grid.
Figure 2: The pipeline of our system. There are three modules in our system: depth estimation module, Feature-metric Bundle Adjustment pose estimation module, and incremental scene implicit representation module. All three modules are jointly optimized during the system operation.
II-C NeRF with Pose Optimization
The original NeRF method needs groundtruth pose information as their input, which is difficult to obtain in real-world environments. They require pre-computed camera parameters obtained from SFM and SLAM algorithms [34, 35, 36]. Recently, many researchers have focused on implicit scene representation without pose prior. iNeRF [37] is the first work focused on pose estimation and scene representation. They estimate the camera pose for novel view images with a pre-trained NeRF model. NeRFmm [38] jointly optimised as learnable parameters with NeRF training, through a photometric reconstruction. BARF [3] proposes a coarse-to-fine Bundle Adjustment method to estimate camera poses with neural 3D representation. Some other methods [39, 40, 41, 42] combine semantic feature with NeRF to improve the scene representation and camera tracking. NoPe-NeRF [9] and LocalRF [10] are the most relevant work to ours. NoPe-NeRF learns scale and shift parameters during training and incorporates undistorted monocular depth prior. However, they often fail in large-scale scene representation due to the limited scene representation capacity. Their pose estimation module are not robust and accurate in large-scale scenes, resulting in increased pose drifts.
@@ III-A Depth Estimation
III-B Feature-metric BA Pose Estimation
In terms of pose estimation, we propose a FBA network. We show the architecture of our network in Fig. 2. Inspired by [46], each image is encoded into high-level features and uses FBA to get an accurate and robust pose estimation. For every consecutive frames $I_{a},I_{b}$, we use a Unet-like encoder-decoder network to extract $D_{l}$ dimensional feature map $\mathbf{F}^{l}\in\mathbb{R}^{W_{l}\times H_{l}\times D_{l}}$ at each level $l\in\{L,\ldots,1\}$. Those multi-channel image features with decreasing resolution can encode richer semantic information and larger spatial information of the image. Our network can also simultaneously estimate the confidence of each pixel to select the robust features for pose estimation. Compared with the original BA method, this learned representation is robust to large illumination or viewpoint changes, which can provide accurate pose estimation in indoor and outdoor environments. Despite poor initial pose estimates, it can also provide meaningful gradients for successful estimation. Those joint learning methods with classical direct alignment [9, 10] operate on the original RGB image, which is not robust to large-scale scenes and long-term changes, and resorts to Gaussian image pyramids, which still largely limits the convergence to frame-to-frame tracking.
Feature Residual: We define the direct feature residual between the two consecutive frames. We optimize the relative pose$\{\mathbf{R},\mathbf{t}\}$ through minimizing the feature residual. For a given feature level $l$ and each 3D point $i$ observed in two consecutive frames, we define a residual:
where $\mathbf{p}_{a}^{i}=\Pi\left(\mathbf{R}\mathbf{P}_{i}+\mathbf{t}\right)_{b}$ is the projection of frame b given its current pose estimate. []delimited-[][$\cdot$][ ⋅ ] is a lookup with sub-pixel interpolation. $F_{a}^{l},F_{b}^{l}$ are the corresponding feature maps of images a and b. The total error over N observations is:
Algorithm 1 Incremental Implicit Scene Representation.
1:  $j\leftarrow 1$; /$\ast$/ ∗ The first scene representation index/$\ast$/∗ /
2:  $p\leftarrow 1$; /$\ast$/ ∗ First camera pose index/$\ast$/∗ /
3:  $q\leftarrow 1$; /$\ast$/ ∗Start with the first frames $\ast$/∗ /
4:  $\{R,T\}_{p..q}\leftarrow 1$; /$\ast$/ ∗ Initialize poses as identity/$\ast$/∗ /
5:  $\theta_{j}\leftarrow InitializeLR()$; /$\ast$/ ∗ Initialize the first local radiance fields/$\ast$/∗ /
6:  while $q<P$ do
7:     while $m=0;m<100$ do
8:        Optimize $(\theta_{j})$; /$\ast$/ ∗ Optimize the first local radiance field with the first frame/$\ast$/∗ /
9:        $m\leftarrow m+1$;
10:     end
11:     while $\{R,t\}_{j}<Bound_{j}$ do
12:        $q\leftarrow q+1$;
13:        $\{R,T\}_{q}\leftarrow\{R,T\}_{q-1}$; /$\ast$/ ∗ Append a pose at the end of the trajectory $\ast$/∗ /
[…]
15:     end
16:     if $q<P$ then
17:        $j\leftarrow j+1$; /$\ast$/ ∗ A new local radiance field index/$\ast$/∗ /
18:        $\theta_{j}\leftarrow InitializeLR()$ /$\ast$/ ∗ Initialize a new local radiance field/$\ast$/∗ /
19:        $t_{j}\leftarrow t_{q}$; /$\ast$/ ∗ Centered around the last pose/$\ast$/∗ /
20:        $p\leftarrow p+n$; /$\ast$/ ∗ Stop considering the frames in last local radiance field/$\ast$/∗ /
21:     end if
22:     Repeat until all frames are registered.
23:  end
@@ III-B Feature-metric BA Pose Estimation
Tanks and Temples [8] Ballroom 20.66 0.50 0.30 24.97 0.71 0.20 25.29 0.89 0.08 26.50 0.85 0.08
Barn 25.28 0.64 0.24 25.77 0.68 0.22 27.89 0.88 0.11 28.59 0.91 0.11
Church 23.17 0.62 0.27 23.60 0.67 0.23 28.07 0.91 0.07 28.82 0.92 0.08
Family 23.04 0.61 0.28 23.77 0.68 0.22 28.60 0.92 0.007 30.48 0.94 0.05
Francis 25.85 0.69 0.28 29.48 0.80 0.18 31.71 0.93 0.11 33.38 0.93 0.10
Horse 24.09 0.72 0.21 27.30 0.83 0.13 29.98 0.94 0.07 30.07 0.93 0.07
Ignatius 21.78 0.47 0.30 23.77 0.61 0.23 26.54 0.86 0.09 27.33 0.87 0.09
Museum 23.58 0.61 0.28 26.39 0.76 0.17 26.35 0.86 0.16 29.43 0.92 0.12
Mean 23.42 0.61 0.27 26.34 0.74 0.19 27.87 0.91 0.09 29.04 0.90 0.09
@@ III-C Incremental Implicit Scene Representation
Existing scene implicit scene representation methods such as [4, 3] can only achieve satisfactory results in small-scale scenes. They are frequently constrained by their limitations in scene representation capabilities when dealing with longer sequences and tend to converge towards local minima during the optimization process. Inspired by [10], we propose an incremental implicit scene representation. The architecture of our network is shown in Fig. 2 and Alg. 1. At the beginning, we will initialize a local neural radiance field and optimize it using the initial few frames as input. As the robot moves forward, we progressively append new frames within the local neural radiance field and iteratively refine the scene representation, pose estimation, and depth estimation. This procedure persists until we reach the boundary of the local neural radiance field. Then, we dynamically allocate new local radiance fields throughout the optimization with frames within a sliding window. This procedure will continue until all the images are registered into the network. So, the entire scene can be represented as multiple local NeRF:
@@
$T_{i}$ $=\exp\left(-\sum_{i-1}^{j}\sigma_{j}\delta_{j}\right)$
Depth Guided Sampling  The estimated depth images provide valuable geometry information which can guide neural point sampling along a ray. We get $N_{strat}$ points for stratified sampling between the near and far planes. Then, $N_{surface}$ points are drawn from the Gaussian distribution determined by the depth prior. When the depth is not known or invalid, we use the rendered depth. Compared to the original methods, our approach allows for more efficient point sampling and enhances the scene representation capability of the network.
Figure 4: Qualitative results on Static Hikes Indoor datasets [10] with long trajectory and span over 1000 images. We present the scene reconstruction results of NoPe-NeRF[9], LocalRF [10], and our method in different time steps (90, 220, 390, 700). Our incremental scene representation method significantly improves the scene reconstruction performance in large-scale scenes.
III-D Incremental Joint Learning
$\mathcal{L}_{d}=\left|\hat{\mathbf{D}}^{*}-\mathbf{D}^{*}\right|$ (10)
where $\hat{D}^{*}(r)$ is the normalized render depth, $D^{*}(r)$ is the normalized estimated depth image from our depth network. We use a pre-trained model to get depth images prior to our depth network. Since the monocular depth is not scale- and shift-invariant, we normalize the depth and estimate the scale and shift of depth:
$t(D)=\frac{1}{N}\sum_{i=1}^{N}(D)$ (11)
$s(D)=\frac{1}{N}\sum_{i=1}^{N}|D-t(D)|$ (12)
We sample M pixels for the frame. Then, we normalize the depth:
$D^{*}=\frac{D-t(D)}{s(D)}$ (13)
We define a photo-metric loss to optimize local radiance fields:
@@ III-D Incremental Joint Learning
This is the forward optical flow, and we also calculate backward optical flow $L_{fb}$ to supervise our implicit scene representation network. The final loss of our network is :
@@
TABLE III: The view synthesis results on Tanks and temples dataset. We report the average PSNR, SSIM, LPIPS results.
[…]
COLMAP NeRFfacto[48] 24.17 0.65 0.194
Mip-NeRF360[47] 27.74 0.77 0.152
NoPe-NeRF[9] 26.43 0.74 0.184
LocalRF[10] 28.04 0.88 0.096
Ours 28.17 0.88 0.095
Joint learning SCNeRF [4] 20.82 0.58 0.227
NeRFmm [38] 22.53 0.64 0.297
BARF [3] 23.42 0.64 0.275
NoPe-NeRF [9] 26.34 0.74 0.194
LocalRF [10] 27.87 0.91 0.094
Ours 29.04 0.90 0.092
IV Experiments
Baselines Methods and Public Datasets. The main baseline methods we compared are BARF [3], NoPe-NeRF [9], and LocalRF [10]. All of them are joint learning methods of camera poses and implicit scene representation. We also compare our method against SCNeRF [4], NeRFmm [38], NeRFfacto [48]. We evaluate our method on the Tanks and Temples dataset [8] and Static Hikes dataset [10]. The Tanks and Temples dataset features slow-moving image sequences spanning a few meters in both indoor and outdoor scenes. We also utilize the Static Hikes indoor dataset, which includes hand-held sequences with longer camera trajectories, testing scalability and pose estimation robustness over ranges typically exceeding 100 meters. We strictly follow the experimental settings of our baseline to partition the training set and the test set.
Implementation Details. We present our implementation details. We select Adam optimizer[50] ($\beta=(0.9,0.999)$) for scene representation and camera tracking optimization. The color loss weighting is $\lambda_{1}=0.25$, $\lambda_{2}=0.1$ for depth and $\lambda_{3}=1.0$ for pose estimation, $\lambda_{4}=1.0$, $\lambda_{5}=1.0$ for optical flow. All training and evaluation experiments are conducted on a single NVIDIA RTX 3090 GPU. Our initial learning rate are $5\times 10^{-3}$ for rotations and $5\times 10^{-4}$ for translations. The resolution of our local radiance fields is $64^{3}$. When the camera moves to the bound, we refine the local radiance fields and the camera pose for 600 iterations per frame and upsamle the resolution to $640^{3}$. After that, we allocate new local radiance fields and repeat this process until all frames are registered.
TABLE IV: Novel view synthesis on Static Hikes dataset [10] and Proprietary dataset. We use PSNR, SSIM, and LPIPS as our metrics.
Methods Static Hikes Indoor [10] Our Dataset